Mr Jimenez has 8 full packages of notebook paper in his cabinet. On his desk there are two 1/4 full packages. There are seven 1/2 full packages on the supply table at the front of the classroom. Each week mr jimenezs classes use enough notebook paper to fill 5 whole packages. How long will it take mr jimenezs classes to use all of the notebook paper he currently has?
Answer by richwmiller(17219)

(8+2*1/4+7*1/2)/5=
(8+.5+3.5)/5=
12/5= 2.5 weeks